OpenStudy (anonymous):

1.which of newtons law describes inertia? 2.which of newtons laws implies that it takes more force to accelerate a heavy object than a light object? 3.what force could cause a moving object to stop? 4.what is the mathematical formula that describes newtons second law?

OpenStudy (anonymous):

Isaac Newton defined inertia as his first law in his Philosophiæ Naturalis Principia Mathematica, which states:[1] The vis insita, or innate force of matter, is a power of resisting by which every body, as much as in it lies, endeavours to preserve its present state, whether it be of rest or of moving uniformly forward in a straight line.

OpenStudy (anonymous):

Newton's Second Law of Motion Newton's Second Law of Motion states that when a force acts on an object, it will cause the object to accelerate. The larger the mass of the object, the greater the force will need to be to cause it to accelerate. This Law may be written as force = mass x acceleration or: F = m * a Another way to state the Second Law is to say it takes more force to move a heavy object than it does to move a light object. Simple, right?

OpenStudy (anonymous):

2. second law F = m * a 3. friction 4. thats given in 2. another form of 2nd law is F = d(mv)dt

OpenStudy (anonymous):

no.1 question sir?

OpenStudy (anonymous):

much more simple

OpenStudy (anonymous):

not sure about that one ti be honest

OpenStudy (anonymous):

thx sir